OpenStudy (anonymous):

what is the value of A -5(a+3)=55

OpenStudy (anonymous):

haseeb96 dont just give him an answer

OpenStudy (solomonzelman):

divide both sides by '-5' subtract 3 from both sides

OpenStudy (anonymous):

Your mad cause he was gonna give it me faster than you.. but would it be -14

OpenStudy (haseeb96):

-5(a+3)=55 -5a-15=55 -5a=55+15 -5a=70 a=-70\5 a=-24 the answer is -24

OpenStudy (anonymous):

no its not^

OpenStudy (haseeb96):

clear it

OpenStudy (anonymous):

heres my best guesses 8 -8 14 -14

OpenStudy (anonymous):

-8 right

OpenStudy (haseeb96):

-14 sorry i forgot it

OpenStudy (anonymous):

so i was right its -14

OpenStudy (anonymous):

-5(a+3)=55 =>a+3=-55/5 =>a=-11-3 =>a=-14

OpenStudy (haseeb96):

instead of 1 i wrote 2

OpenStudy (solomonzelman):

@Haseeb96 and @SuyashGupta, guide the asker, don't preform his/her steps and his work! @jrc5101520, ty for your cooperation! @Haseeb96, the goal is to make the asker understand not give him/her work with answer.

OpenStudy (anonymous):

@Haseeb96 oo... dats da worst mistake i do in exams.. lol

OpenStudy (anonymous):

@SolomonZelman okay bro.. sorry..

OpenStudy (solomonzelman):

@SuyashGupta, it's OK, just be cooperative in the future! you are new to the website, and now you know! \[(•_‿•) \]